Accounting 305

Chapters 2 & 3 Homework

  1. Hometown Construction Corp. entered into a long-term contract to build an office building for Outsider Corp. at a fixed contract price of $27.5 million on 6/15/2016. Construction of the building is completed on 12/15/2018. Hometown Construction expects the building to be completed in two and a half years and uses the Percentage of Completion Method (cost to cost approach) to account for the contract. Cost information related to the contract are as follows:

2016

2017

2018

Actual Construction costs incurred during year

4,500,000

12,075,000

9,425,000

Estimated costs to complete at 12/31

20,500,000

8,925,000

0

A.         Determine the amount of Gross Profit that would be recognized each year under Percentage of Completion.

B.         What amount of Gross Profit would be recognized in 2017 if the estimated costs to complete were $11,000,000?

C.        What would be the amount of Gross Profit recognized each year if the completed Contract Method were used instead?
Answer #1
A.) % Completion of 2016 18.00% =4500000/(4500000+20500000)
Revenue to be recognized in 2016        4,950,000 =27500000*18%
Less: Cost incurred in 2016        4,500,000
Gross profit of 2016 $ 450,000
% Completion of 2017 65.00% =(4500000+12075000)/(4500000+12075000+8925000)
Cumulative revenue at end of 2017     17,875,000 =27500000*65%
Less: Revenue recognized in 2016        4,950,000
Revenue to be recognized in 2017    12,925,000
Gross profit of 2017 $ 850,000 =12925000-12075000
Revenue to be recognized in 2018        9,625,000 =27500000-4950000-12925000
Less: Cost Incurred in 2018        9,425,000
Gross profit of 2018 $ 200,000
B.) % Completion of 2017 60.11% =(4500000+12075000)/(4500000+12075000+11000000)
Cumulative revenue at end of 2017     16,530,250 =27500000*60.11%
Less: Revenue recognized in 2016        4,950,000
Revenue to be recognized in 2017    11,580,250
Gross profit(loss) of 2017        - $494,750 =11580250-12075000
C.) Gross profit of 2016 $ 0
Gross profit of 2017 $ 0
Gross profit of 2018 $ 1,500,000